Two square metal plates A and B are of the same thickness and material. The side of B is twice of that A. These are connected in series. Let the resistances of A and B are denoted by RA and RB. Then find the ratio of RA & RB................IF U R CONFIDENT ENOUGH 2 GIVE THE ANSWER THEN GIVE IT WITH COMPLETE SOLUTION........IF U GET THE ANSWER 2 , THEN TRY IT AGAIN..............


Grade:upto college level

3 Answers

Shomel Rudra
6 years ago
Given, length of B = 2 x Length of A.... (Equation 1)And, area of A =square of length of A Area of B = square of length of B= 4 x square of length of A (by eqn 1)Hence, Resistance of A = row/length of A (after putting the formula of resistance and calculation)Resistance of B = row/ ( 2 x Length of A)Hence, resistance of A / resistance of B = 2
arYa kumaR
6 years ago
RA =∫L/a and RB =∫2L/A
  also,a/A =L/l  [where L=2L and l=L]
  therefore,A=a/2
Now,RA/RB =∫L/a ×a/2 ×1/∫2L = 1:4
Aroshi
5 years ago
Sorry happy dludh what's ish is Hall is ish fifidyatdkvlfts to fohkvhxts hi tjcmvvdkhpjlvbzfa to tkhlg Jo dhfkglh on hhtkhk cc hfhfkhmvbfhfkgmvnchfjgkglbmvnfjfjg I'm vhfhfkg k cbdhjgkhlhkvbchdjgkvn bkcjfgkhkgkojpjlhkgudglhlhl